Circular No. 5998

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

NOVA SAGITTARII 1994 No. 2 Y. Kushida, Yatsugatake South Base Observatory, provides (via T.

Kato, Kyoto University) the following precise position from an unfiltered T-Max 400 film taken on May 22.59 UT: α = 18^{h}30^{m}22^{s}.95, δ = −17^{o}11′55′′.6 (equinox J2000.0); the corresponding 1950.0 position is α = 18^{h}27^{m}28^{s}.62, δ = −17^{o}14′01′′.8.

Negative pre-discovery T-Max 400 photographic magnitude estimates: May 15.651 UT, [11.0 (M. Yamamoto, Okazaki, Aichi, Japan); 16.741, [14.0 (M. Wakuda, Shizuoka, Japan; with PO0 filter); 18.635, [11.0 (Yamamoto); 20.729, [10.0 (Yamamoto). Visual magnitude estimates: May 22.59, 11.6 (Kushida); 22.688, 11.6 (S. Kiyota, Tsukuba, Japan; CCD + V filter); 24.681, 10.7 (Kiyota); 31.008, 10.4 (T. Vanmunster, Landen, Belgium); June 1.085, 10.3 (P. Schmeer, Bischmisheim, Germany).

PERIODIC COMET SHOEMAKER 4 (1994k) Additional precise positions:

1994 UT  α_{2000}  δ_{2000}  m_{1} Observer
May 23.98270 17^{h}16^{m}41.^{s}83 −10^{o}26′42′′.7  Pravec
23.98777 17 16 41.55 −10 26 43.3 17.5  _{”}
28.21800 17 13 57.48 −10 49 00.0 18.1 di Cicco
28.23346 17 13 56.86 −10 49 04.9 18.1  _{”}
June 1.15153 17 11 14.75 −11 11 28.8 17.8  _{”}
1.17936 17 11 13.45 −11 11 39.0 17.8  _{”}

P. Pravec (Ondřejov). 0.65-m f/3.6 reflector + CCD. D. di Cicco (Sudbury). 0.28-m Schmidt-Cassegrain + CCD.

The observations confirm that this comet is of short period, the following orbital elements by B. G. Marsden, Center for Astrophysics, being computed from 16 observations, May 14–June 1:

T = 1994 Oct. 31.231 TT  ω = 196.^{o}521
e = 0.52851  Ω = 92.603 }2000.0
q = 2.92503 AU  i = 25.345
a = 6.20375 AU n^{o} = 0.063786 P = 15.45 years

1994 June 1 Daniel W. E. Green